# SMC++: Masked Learning of Unsupervised Video Semantic Compression

### v1

@@ 4.4 Ablation Studies and Analysis for SMC Model
TABLE V: Impact of fine-tuning (FT) downstream model to different coding methods on UCF101-TSM (Top1) evaluation setting at the 0.04bpp bitrate level.
VVC VVC+BasicVSR++ SMC
Before FT(%) 76.97 82.46 86.46
After FT(%) 82.35(+5.38) 82.92(+0.46) 89.83(+3.37)
We also fine-tune the official UCF101-TSM model using videos compressed by the above variant methods. As shown in Table V, the performance of vanilla VVC after fine-tuning (FT) nearly catches up with that of BasicVSR++ (82.46%), suggesting that the video enhancement procedure mainly narrows the domain gap between the evaluation and training video data input to TSM model, and bring no new information. In contrast, our full SMC framework further improves the accuracy from 86.46% (already higher than BasicVSR++) to 89.83% at 0.04bpp level. This proves that our method consistently benefits the downstream task by compensating new semantic information, whether fine-tuning the downstream model or not.
TABLE VI: Impact of different self-supervised learning objectives to SMC performance, on action recognition and MOT tasks. Our masked learning paradigm performs much better, especially at the fine-grained MOT task.
@@
As shown in Table VI, our MAE-powered approach consistently outperforms the contrastive learning-based method MoCoV3 on two tasks of different granularities, i.e., action recognition and MOT, by 1.28% Top1 accuracy and 3.22% MOTA, respectively. This clearly proves the semantics learned by MAE is more generalizable than that learned by contrastive learning, as well as more fine-grained.
@@
Analysis on Masked Learning Objective. We first analyze the correlation between the masked learning loss $\mathcal{L}_{MAE}$ and the semantic fidelity of the compressed video, which is defined as,
$\operatorname{Sem-FID}(\hat{X})=-\log_{10}(\cos(\mathcal{F}(X),\mathcal{F}(\hat{X}))),$ (7)
where $\mathcal{F}$ denotes the video feature extractor network, $\cos$ denotes the cosine distance, i.e., $\cos(\mathbf{a},\mathbf{b})=1-\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\cdot\|\mathbf{b}\|}$. The larger $\operatorname{Sem-FID}$ value, the better the semantic similarity of the compressed video to the original video. As shown in Figure 9, the masked learning objective $\mathcal{L}_{MAE}$ consistently serves as an effective semantic learning proxy for different downstream models such as TSM and TimeSformer.
Additionally, we visualize the videos reconstructed from the masked compressed videos under different bitrate levels. As shown in Figure 10, with more bitcost (0.0168$\rightarrow$0.0189bpp), the masked learning loss $\mathcal{L}_{MAE}$ is reduced (0.0727$\rightarrow$0.0703), resulting in better semantic fidelity, as quantified by larger $Sem-FID$ metric (1.078$\rightarrow$1.339). Despite a marginal reduction in $\mathcal{L}_{MAE}$, there is a remarkable enhancement in semantic completeness, as evident in the areas marked by the red arrows. The analyses above consistently prove that the masked learning objective $\mathcal{L}_{MAE}$ is a good surrogate objective for facilitating the compressed video to maintain the semantic information.
@@
Figure 9: Correlation analysis between $\mathcal{L}_{MAE}$ and semantic fidelity of the reconstructed videos by SMC.
Figure 10: Visualization of the MAE reconstructed videos at various bitrate levels. A lower MAE loss ($4th$ row) leads to improved semantic appearance, particularly in the eyes and mouth regions, compared to a higher loss ($3rd$ row).
Further, we study the influence of the masking ratio on the semantic compression performance. As shown in Table VII, the 90% masking ratio achieves the optimal performance, and setting the masking ratio to 75% and 95% achieves slightly inferior performances. The 50% masking ratio shows a remarkable performance drop, as the textures can be simply completed by extending local textures, instead of reasoning high-level scene and structure semantics.
TABLE VII: Impact of different masking ratios to SMC performance, on the UCF101-TSM setting at 0.02bpp.
@@
Top1(%) 73.85 75.62 76.48 76.08
@@
TABLE VIII: Comparison of different non-semantic suppression (NSS) strategies on UCF101-TSM setting. $Res$ (bpp) denotes the bitcost of the residual semantic stream of SMC. The QP of the base codec VVC is set to 51.
@@
To study the necessity of this design, we first train a variant model SMC woNSS by removing the $\mathcal{L}_{NSS}$ item from the loss function. As shown in Table VIII, the bitcost of SMC woNSS is 2.6$\times$ larger than the full SMC model, because the low-level pixel information is back-propagated to the coding system without any selection. Moreover, this non-semantic information may be noises to the downstream tasks, i.e., the action recognition accuracy is dropped from 72.96% to 69.57%. Then, we train a variant model SMC woSem by using a plain learnable variable as the condition of the GMM distribution, instead of the semantic feature $S$. Both the compression efficiency and the recognition accuracy of SMC woSem are superior to SMC woNSS, but still inferior to our SMC model. This implies that explicitly regularizing the information entropy of the MAE token space is beneficial to a coding system, and using learned semantics as the guidance further improves this idea. Finally, we use the vector quantization (VQ) [147] codebook to discretize the token space of MAE, and the resultant SMC VQ model has similar performance to SMC woSem, indicating that the idea of information suppression is important for a compression system, instead of its concrete implementation.
Hyper-parameter Sensitivity. Setting the semantic loss item weight $\alpha$ to the values in the range [1,10] gives similar results. Setting the NSS loss item weight $\beta$ to 0.1 achieve the best results, while a smaller value of 0.01 cannot effectively suppress the non-semantic information and a larger value of 1 makes the semantic learning item hard to optimize. Setting the GMM component number K to values in the range [3,9] gives similar results, while smaller K is slightly worse.
@@ 4.5 Analysis on the Improved SMC++ Model
Figure 13: Ablation study on the SMC++ framework: (1) SMCadv shares the same details as the original SMC, but with extended training iterations to align with SMC++ for a fair comparison. (2) The SMCadv+Motion model is trained with both masked appearance and motion modeling objectives. (3) The SMC++ model further improves SMCadv+Motion model with Blue-Tr.
@@
As shown in Figure 14 (a), introducing motion prediction branch to the MAE learning objective always leads to the performance improvement. Compared to the computationally expensive feature-level matching operations in computing optical flow and correlation map, obtaining RGB difference map is more efficient. However, it demonstrates the strongest result, over 1% average Top1 accuracy improvement over the baseline SMCadv model, obviously superior to the gains 0.15% and 0.6% that are obtained by optical flow and feature correlation. This is not surprising, as using frame difference as motion representation has effectively enhanced various video analysis tasks [151][84][152].
@@
Ablation Study on Blue-Tr. As shown in Figure 15 (a), all variant modules of Blue-Tr outperform the CNN-based basic compression module Basic-CM. This proves the advantage of Transformer over CNN in compressing highly abstract semantic information.
@@
Figure 15: (a) Ablation study of the proposed Blueprint-guide compression Transformer (Blue-Tr). “woBlueprint” denotes the blueprint-based alignment design is removed. “woDomainAlign” represents the domain alignment operation in Feature Aligner is excluded, and only the motion alignment is preserved. All models are trained with both masked-appearance and masked-motion modeling objectives for a fair comparison. (b) Comparison of the adopted decomposed Transformer and the vanilla full Transformer.
Furthermore, we note that removing the blueprint-based alignment, the resultant Blue-Tr woBlueprint model depicts a notable decrease in Top1 accuracy, i.e., about 5% drop at a 0.04bpp bitrate. The reason is that, without explicit feature alignment, the Transformer still struggles to capture redundancies among various contextual features.
Moreover, to assess the impact of domain correction and motion compensation, we removed the Domain-Net from the Feature Aligner. The resulting model, Blue-Tr woDomainAlign, performs obviously worse than the full Blue-Tr, particularly at low bitrates. For instance, at 0.025bpp, there’s a performance dip of over 3%. This decline is attributed to the domain gap between the features of the lossy video (compressed by the base VVC codec) and our semantic features, which is further exacerbated by the severe semantic distortion of VVC in low-bitrate conditions.
We also investigate the effectiveness of the decomposed Transformer used in Blue-Tr. Specifically, a point-wise Transformer first removes redundancy among feature types, followed by a spatial Transformer to eliminate spatial redundancy. Figure 15 (b) demonstrates that our decomposition strategy enhances speed by about 2x compared to the standard Transformer (10.9ms vs. 20.2ms), while performance remains similar. Notably, our method slightly outperforms the standard Transformer at lower bitrate levels, achieving 47.75% vs. 47.23% at 0.03bpp, potentially because decomposition mitigate the feature ambiguity.
TABLE IX: Comparison of different Blueprint feature dimensions on the Diving48-TSM setting at 0.04bpp bitrate level.
@@
Top1 (%) 53.11 54.02 54.58 54.26 54.13
Finally, we explore how the blueprint semantic feature’s dimension impacts system performance. As shown in Table IX, smaller dimensions of 32 and 64 yield poorer results due to insufficient information for effective feature alignment. Conversely, larger dimensions like 256 and 512, while not causing drastic performance declines, still lead to a noticeable drop. This is because the bitcost savings from blueprint guidance can’t offset the cost of transmitting the blueprint feature itself.
@@ 4.6 Model Complexity and Running Time
As Table X demonstrates, our basic SMC model includes fewer parameters than the recent learnable codec DCVC-FM, but more parameters than the previous semantic coding approach VCS. VCS utilizes complex dynamic convolution in its encoder to enhance feature representation, striking a balance between parameter count and performance, albeit at the expense of a longer encoding time (1023ms vs. SMC’s 952ms). Furthermore, in the VCS, the semantic extraction and compression modules are tightly decoupled within the encoder network. In contrast, our SMC model separately employs a simple standard Resnet18 for semantic extraction, while isolating the semantic compression process into a distinct module. SMC’s simplicity and modular design make it a flexible and easily-extendable baseline.
TABLE X: Model complexity and running time of different approaches. denotes the method adopting the VVEnC software as the base coding layer, where the coding time consumed by VVEnC is incorporated. All evaluations are performed on a machine equipped with a single Nvidia 2080Ti GPU and Intel(R) Xeon(R) Silver 4116 CPU @ 2.10GHz. The running time was determined by encoding/decoding 100 frames of 1080p resolution and subsequently calculating the average time required for a single frame.
Method Parameters Encoding Time Decoding time
VVEnc15.0 - 652ms 142ms
DCVC-FM 59.63M 833ms 654ms
VCS 28.60M 1023ms 198ms
SMC 48.23M 952ms 235ms
SMC++ 96.18M 977ms 253ms
Our SMC++ model further introduces Transformer blocks, increasing the parameter count compared to SMC. However, the execution times are not obviously affected. This is because the new compression module Blue-Tr in SMC++, operating on significantly downsampled feature maps (32$\times$), is extremely efficient.

### v2

@@ 4.3 Ablation Studies and Analysis for SMC Model
TABLE V: Impact of different self-supervised learning objectives to SMC performance, on action recognition and MOT tasks. Our masked learning paradigm performs much better, especially at the fine-grained MOT task.
@@
As shown in Table V, our MAE-powered approach consistently outperforms the contrastive learning-based method MoCoV3 on two tasks of different granularities, i.e., action recognition and MOT, by 1.28% Top1 accuracy and 3.22% MOTA, respectively. This clearly proves the semantics learned by MAE is more generalizable than that learned by contrastive learning, as well as more fine-grained.
@@
Ablation on Visual Quality Loss Terms. Remembering that we ensure the reconstructed video photo-realistic, by using the combined LPIPS plus GAN terms in VQGAN. We conduct more studies to study if the two terms are critical to our problem. As shown in Figure 11, when the GAN loss is removed, the performance degrades by approximately 5% at 0.02 bpp. In contrast, removing the LPIPS loss severely degrades the performance, with an accuracy drop of about 15% at 0.02 bpp. This degradation arises because the local structural features within each patch are no longer preserved, whereas our MAE loss only retains the global interactions among different features. In summary, both the two loss terms are critical for a high-performance video semantic compression system. Furthermore, we explored other weighting strategies for the GAN loss term, including the adaptive weighting approach employed in VQGAN [100]. As shown in Figure 11 (b), these strategies did not result in obvious performance improvements.
@@
  (a)    (b)
Figure 11: (a) Ablation study on the LPIPS+GAN terms for SMC. “w/o” denotes removing the term. (b) Impact of $L_{GAN}$ weight to SMC. “Ada” denotes adaptive weight.
Analysis on Masked Learning Objective. First, we demonstrate that the lower MAE loss is positively related to the video semantic completeness, as demonstrated in Figure 1 and 2 of supplementary material.
Further, we study the influence of the masking ratio on the semantic compression performance. As shown in Table VI, the 90% masking ratio achieves the optimal performance, and setting the masking ratio to 75% and 95% achieves slightly inferior performances. The 50% masking ratio shows a remarkable performance drop, as the textures can be simply completed by extending local textures, instead of reasoning high-level scene and structure semantics.
TABLE VI: Impact of different masking ratios to SMC performance, on the UCF101-TSM setting at 0.02bpp.
@@
Top1 (%)
73.85 75.62 76.48 76.08
TABLE VII: Comparison of different MAE prediction targets for SMC, on UCF101-TSM setting.
Method UCF101-TSM Top1 (%)
@ 0.02bpp @ 0.04bpp @ 0.06bpp @ 0.07bpp
Raw pixel
76.48 86.46 89.11 89.63
DINOv2 feature
77.00 86.48 89.51 89.60
Finally, we explore using features extracted from pre-trained neural networks as prediction targets for the MAE task, instead of raw pixels. Specifically, we employ the pre-trained DINOv2 [141] model to extract features, which serve as the prediction target during masked learning.
As shown in Table VII, the DINOv2-based approach achieves better results at very low bitrates due to its compact feature representation. However, at higher bitrates, it underperforms the pixel-prediction approach. For example, at 0.06 bpp, the DINOv2-based model performs 0.60% worse. Additionally, the feature extraction process increases training time by 32%. These results indicate that while feature-based targets are beneficial for extremely low bitrates, pixel-based targets remain stable across all bitrate levels.
@@
TABLE VIII: Comparison of different non-semantic suppression (NSS) strategies on UCF101-TSM setting. ResRes (bpp) denotes the bitcost of the residual semantic stream of SMC. The QP of the base codec VVC is set to 51.
@@
TABLE IX: Impact of $\beta$ to SMC, on UCF101-TSM@0.02bpp.
$\beta$ 0.01 0.05 0.1 0.5 1
Top1(%) 72.94 75.21 76.48 76.23 75.93
TABLE X: Impact of $K$ to SMC, on UCF101-TSM@0.02bpp.
$K$ 1 3 5 7 9 11 15
Top1 (%) 76.13 76.42 76.48 76.47 76.42 76.23 75.92
[…]
Hyper-parameter Sensitivity. Setting the semantic loss item weight $\alpha$ to the values in the range [1,10] gives similar results. Then, we investigate the impact of $\beta$, which balances the Non-Semantics Suppressing term and the MAE term. As shown in Table IX, increasing $\beta$ from 0.01 to 0.1 leads to consistent performance improvements, due to the increased semantic separability. However, when $\beta$ becomes too large, the representation capability of the video token features in MAE becomes overly constrained, resulting in degraded performance. Regarding the GMM component number $K$, setting it within the range $[3,9]$ yields comparable results, with $K=5$ empirically achieving optimal performance, as demonstrated in Table X. A small value such as $K=1$ results in an overly simplistic mixture distribution that inadequately captures the semantic features, while excessively large values such as $K=15$ create distributions that are too complex for effective optimization.
@@ 4.4 Analysis on the Improved SMC++ Model
Refer to caption Refer to caption Refer to caption
Figure 14: Ablation study of SMC++ framework on action recognition, MOT and VOS tasks. ‘SMCadv’ shares the same details as the original SMC, but with extended training iterations to align with SMC++ for a fair comparison. ‘SMCadv+Motion’ model is trained with both masked appearance and motion modeling objectives. SMC++ model further improves ‘SMCadv+Motion’ model with Blue-Tr.
@@
Furthermore, we conduct ablation studies on MOT and VOS tasks, to comprehensively validate the effectiveness of the masked motion modeling scheme and the Blue-Tr module. As shown in Figure 14, both the masked motion objective and the Blue-Tr module contribute to the superior performance on the VOS and MOT tasks. For instance, in the VOS task, adding the masked motion objective results in the SMCadv+Motion model achieving 79.76% in $\mathcal{J\&F}$ at 0.02bpp, which is 0.46% higher than the 79.30% attained by SMCadv. Furthermore, by integrating the transformer-based Blue-Tr module, our final SMC++ model reaches 81.06% in $\mathcal{J\&F}$ at 0.02bpp, marking a further improvement of 1.30% over the SMCadv+Motion model.
These results demonstrate that the two new modules introduced in SMC++ are always effective, across tasks of varying granularity, from classification to segmentation.
@@
As shown in Figure 15 (a), introducing motion prediction branch to the MAE learning objective always leads to the performance improvement. Compared to the computationally expensive feature-level matching operations in computing optical flow and correlation map, obtaining RGB difference map is more efficient. However, it demonstrates the strongest result, over 1% average Top1 accuracy improvement over the baseline SMCadv model, obviously superior to the gains 0.15% and 0.6% that are obtained by optical flow and feature correlation. This is not surprising, as using frame difference as motion representation has effectively enhanced various video analysis tasks [85].
@@
Finally, we investigate the balancing weight between masked pixel and motion prediction targets, by introducing the tunable hyperparameter $\gamma$ instead of the fixed 0.5, i.e.,
$\mathcal{L}_{MAE}=(1-\gamma)\mathcal{L}_{MAE-pixel}+\gamma\mathcal{L}_{MAE-motion}.$ (25)
As shown in Table XI, setting $\gamma=0.5$ consistently yields optimal or near-optimal performance across various bitrates. For example, at 0.06 bpp, the Top1 accuracy reaches 59.54%, outperforming both $\gamma=0$ (58.87%) and $\gamma=1$ (58.91%). At lower bitrates, such as 0.025 bpp, the benefit of higher $\gamma$ values becomes evident, with $\gamma=0.5$ achieving 35.77% and $\gamma=1$ achieving 36.13%. This is likely because motion information is sparser and can be more efficiently compressed at lower bitrates.
TABLE XI: Comparison of different tunable balancing weights ($\gamma$) between the pixel-based and motion-based MAE learning target in the SMCadv+Motion model.
$\gamma$ Diving48-TSM Top1 Accuracy (%)
@0.025bpp @0.04bpp @0.06bpp @0.08bpp
0 33.40 50.10 58.87 62.83
0.25 34.44 50.28 58.90 62.88
0.5 35.77 50.59 59.54 63.45
0.75 35.97 50.50 59.29 63.13
1 36.13 50.37 58.91 62.73
Ablation Study on Blue-Tr. As shown in Figure 16 (a), all variant modules of Blue-Tr outperform the CNN-based basic compression module Basic-CM. This proves the advantage of Transformer over CNN in compressing highly abstract semantic information. Furthermore, we note that removing the blueprint-based alignment, the resultant ‘Blue-Tr woBlueprint’ model depicts a notable decrease in Top1 accuracy, i.e., about 5% drop at a 0.04bpp bitrate. The reason is that, without explicit feature alignment, the Transformer still struggles to capture redundancies among various contextual features.
@@
Figure 16: (a) Ablation study of Blueprint-guide compression Transformer (Blue-Tr). (b) Comparison of the adopted decomposed Transformer and the vanilla full Transformer.
Moreover, to assess the impact of domain correction and motion compensation, we removed the Domain-Net from the Feature Aligner. The resulting model ‘Blue-Tr woDomainAlign’ performs obviously worse than the full Blue-Tr, particularly at low bitrates. For instance, at 0.025bpp, there’s a performance dip of over 3%. This decline is attributed to the domain gap between the features of the lossy video (compressed by the base VVC codec) and our semantic features, which is further exacerbated by the severe semantic distortion of VVC in low-bitrate conditions.
[…]
We explore replacing full attention with Linear attention [146] to reduce attention complexity from quadratic to linear. As shown in Figure 6 of the supplementary material, the two methods perform similarly at higher bitrates (>>0.06 bpp). However, at lower bitrates such as 0.03bpp, full attention substantially outperforms Linear attention by about 5% accuracy, due to the latter’s limited capability in global context modeling, which is critical for semantic tasks.
Moreover, we explore how the blueprint semantic feature’s dimension impacts system performance. As shown in Table XII, smaller dimensions of 32 and 64 yield poorer results due to insufficient information for effective feature alignment. Conversely, larger dimensions like 256 and 512, while not causing drastic performance declines, still lead to a noticeable drop. This is because the bitcost savings from blueprint guidance can’t offset the cost of transmitting the blueprint feature itself.
TABLE XII: Comparison of different Blueprint feature dimensions, on the Diving48-TSM setting at 0.04bpp bitrate level.
@@
Top1 (%)
53.11 54.02 54.58 54.26 54.13
Previous works [147, 148] have demonstrated that multi-frame reference improves coding performance. To evaluate whether our Blur-Tr module can effectively exploit multi-frame historical information, we incorporate long-term history by aligning multiple previous frames to the current frame using the Feature Aligner in the Blur-Tr module. The aligned features are fused through a residual block, providing a richer coding context compared to using a single previous frame. As shown in Table XIII, incorporating additional previous frames consistently enhances performance, aligning with prior findings [147]. Future work will focus on developing more advanced multi-frame fusion strategies tailored to video semantics, further improving video semantic compression.
TABLE XIII: Comparison of multiple previous frames as the coding context for SMC++, on Diving48-TSM at 0.04bpp.
Previous Frames 1 2 3 4 5
Top1 (%)
54.58 54.83 55.04 55.07 55.12
Finally, we introduce SMC++lite to alleviate Blue-Tr’s increased parameters and computational costs, by using a bottleneck design in the Blueprint-guided compression Transformer. It applies channel reduction before self-attention and expansion afterward. This design cuts computational complexity while maintaining modeling capability. Table XIV shows SMC++lite significantly outperforms SMC (7.28% higher accuracy at 0.02bpp), while maintaining parameter efficiency comparable to SMC. Performance nearly matches the full SMC++ model (only 0.4% lower at 0.02bpp). This confirms our gains come from the novel architecture of Blue-Tr, not just from more parameters.
TABLE XIV: Performance comparison between SMC, SMC++lite, and SMC++ models.
Method Parameters UCF101-TSM Top1 Accuracy (%)
@0.02bpp @0.04bpp @0.06bpp
SMC 48.23M 76.48 86.46 89.11
SMC++lite 52.46M 83.76 88.83 90.81
SMC++ 96.18M 84.16 89.17 90.83
More Advanced Post-Enhancement Strategy. Both SMC and SMC++ adopt the GAN loss to ensure the photo-realism of decoded videos. Recently, stable diffusion [149] has demonstrated advanced image generation capabilities. Consequently, we incorporate the GDP model [150], which leverages the rich generative priors of pre-trained diffusion models [149], to enhance image quality.
As shown in Table XV, the GDP-based enhancement substantially reduces FID (e.g., 27.92 at 0.02 bpp, compared to 76.12 without enhancement), indicating notable improvements in image quality. However, the accuracy improvement is marginal (e.g., 84.56% vs. 84.16%), suggesting that, for semantic compression task, preserving semantic information is more critical than enhancing visual quality.
TABLE XV: Comparison of different approaches, in terms of Top1 Accuracy/Fréchet Inception Distance (FID) [151].
Method UCF101-TSM Top1 (%)/FID
@0.02bpp @0.04bpp @0.06bpp
SMC++ 84.16/76.12 89.17/58.24 90.83/36.53
SMC++(GDP) 84.56/27.92 89.32/26.81 90.89/25.04
Robustness Evaluation. As shown in Figure 7 of the supplementary material, we evaluate SMC++ under varying noise levels (5%, 10%, 20%, and 50%) on the Diving48-TSM setting. As noise intensity increases, accuracy naturally decreases, but the drop remains moderate, e.g.,  2% with 10% noise. Notably, SMC++ demonstrates robustness even at 20% noise levels, likely due to its ability to filter out noise by concentrating on task-relevant features.
@@ 4.5 Model Complexity and Running Time
As Table XVI demonstrates, our basic SMC model includes fewer parameters than the recent learnable codec DCVC-FM, yet more parameters than the previous semantic coding approach VCS. VCS achieves strong feature representation with fewer parameters by employing complex dynamic convolution in its encoder, but this comes at the cost of longer encoding time (1023ms compared to SMC’s 952ms). Furthermore, in the VCS, the semantic extraction and compression modules are tightly decoupled within the encoder network. In contrast, our SMC model separately employs a simple standard Resnet18 for semantic extraction, while isolating the semantic compression process into a distinct module. SMC’s simplicity and modular design make it a flexible and easily-extendable baseline.
[…]
Finally, we provide the complexity and cost of SMC++*, which supports decoding pixel-wise details by appending a lightweight detail decoding network DR-Net to the SMC++ decoder. Compared to SMC++, it only introduces 2.19M parameters while achieving a decoding speed of 296ms, which remains much faster than DCVC-FM (654ms).
TABLE XVI: Model complexity and running time of different approaches. denotes the method adopting the VVEnC software as the base coding layer, where the coding time consumed by VVEnC is incorporated. SMC++* refers to the SMC++ model augmented with DR-Net and further fine-tuned for the PSNR metric. All evaluations are performed on a machine equipped with a single Nvidia 2080Ti GPU and Intel(R) Xeon(R) Silver 4116 CPU @ 2.10GHz. The running time was determined by encoding/decoding 100 frames of 1080p resolution and subsequently calculating the average time required for a single frame.
Method Parameters
Encoding
Time
Decoding
Time
Decoding
FLOPS
VVEnc15.0 - 652ms 142ms -
DCVC-FM 59.63M 833ms 654ms 1651G
VCS 28.60M 1023ms 198ms 147G
SMC 48.23M 952ms 235ms 199G
SMC++ 96.18M 977ms 253ms 307G
SMC++* 98.37M 977ms 296ms 368G